OpenStudy (anonymous):

Which of these is the closes to the value of 1kcal/mole? a) 2/3 * 10^-2 Joules b) 2/3 joules c) 2 * 10^-20 joules d) 3 * 10^-10 Joules.

OpenStudy (espex):

\[1 \frac{kcal}{mol}= 4.184 \frac{kJ}{mol}\]\[1\frac{kJ}{mol}=0.239 \frac{kcal}{mol}\]